# On Sum-Free Functions

### v1

@@ Abstract.
A function from $\mathbb{F}_{2^{n}}$ to $\mathbb{F}_{2^{n}}$ is said to be $k$th order sum-free if the sum of its values over each $k$-dimensional $\mathbb{F}_{2}$-affine subspace of $\mathbb{F}_{2^{n}}$ is nonzero. This notion was recently introduced by C. Carlet as, among other things, a generalization of APN functions. At the center of this new topic is a conjecture about the sum-freedom of the multiplicative inverse function $f_{\text{\rm inv}}(x)=x^{-1}$ (with $0^{-1}$ defined to be 00). It is known that $f_{\text{\rm inv}}$ is 2nd order (equivalently, $(n-2)$th order) sum-free if and only if $n$ is odd, and it is conjectured that for $33\leq k\leq n-$ 3, $f_{\text{\rm inv}}$ is never $k$th order sum-free. The conjecture has been confirmed for even $n$ but remains open for odd $n$. In the present paper, we show that the conjecture holds under each of the following conditions: (1) $n=13$; (2) $3\mid n$; (3) $5\mid n$; (4) the smallest prime divisor $l$ of $n$ satisfies $(l-1)(l+2)\leq(n+1)/2$. We also determine the “right” $q$-ary generalization of the binary multiplicative inverse function $f_{\text{\rm inv}}$ in the context of sum-freedom. This $q$-ary generalization not only maintains most results for its binary version, but also exhibits some extraordinary phenomena that are not observed in the binary case.
@@ 1. Introduction
In a recent paper [4], Carlet introduced the notion of sum-free functions. A function $f:\mathbb{F}_{2^{n}}\to\mathbb{F}_{2^{n}}$ is said to be $k$th order sum-free if $\sum_{x\in A}f(x)\neq 0$ for all $k$-dimensional affine subspaces $A$ of $\mathbb{F}_{2^{n}}$. (Unless specified otherwise, subspaces and affine sunspaces of $\mathbb{F}_{2^{n}}$ are over $\mathbb{F}_{2}$.) This definition does not involve the multiplicative structure of $\mathbb{F}_{2^{n}}$, hence sum-freedom can be defined for functions from $\mathbb{F}_{2}^{n}$ to $\mathbb{F}_{2}^{n}$. However, the multiplicative structure of $\mathbb{F}_{2^{n}}$ plays an important role in the constructions and analysis of such functions. As described in [4], the motivation for the notion of sum-free functions was two-fold: First, sum-free functions are a natural generalization of almost perfect nonlinear (APN) functions; the latter have been extensively studied in cryptography for their resistance against differential attacks [1, 3, 6, 16, 17, 18]. APN functions are precisely 2nd order sum-free functions [9]. Second, sum-freedom addresses the vulnerability of block ciphers to integral attacks which exploit the predicability of the sums of values of S-boxes on affine subspaces [12, 19]. Moreover, from a theoretic point of view, sum-freedom is a notion with rich mathematical content; beneath its simple appearance, there are difficult and enticing questions that require sophisticated mathematical tools.
@@
Let $f_{\text{\rm inv}}$ denote the multiplicative inverse function of $\mathbb{F}_{2^{n}}$, that is, $f_{\text{\rm inv}}(x)=x^{-1}$ for $x\in\mathbb{F}_{2^{n}}^{*}$ and $f_{\text{\rm inv}}(0)=0$. (For $n\geq 2$, we have $f_{\text{\rm inv}}(x)=x^{2^{n}-2}$ for all $x\in\mathbb{F}_{2^{n}}$.) The sum-freedom of $f_{\text{\rm inv}}$ has been the focus of several recent papers [4, 5, 7]. It is well known that $f_{\text{\rm inv}}$ is APN (i.e., 2nd order sum-free) if and only if $n$ is odd [17]. Carlet [4] determined that $\sum_{x\in A}1/x\neq 0$ for all affine subspaces $A$ of $\mathbb{F}_{2^{n}}$ not containing 00. Hence $f_{\text{\rm inv}}$ is $k$th order sum-free if and only if $\sum_{0\neq u\in E}1/u\neq 0$ for all $k$-dimensional subspaces $E$ of $\mathbb{F}_{2^{n}}$. It is trivial that $f_{\text{\rm inv}}$ is $1$st order sum-free and it is known from [5] that for $11\leq k\leq n-$ 1, $f_{\text{\rm inv}}$ is $k$th order sum-free if and only if it is $(n-k)$th order sum-free. Therefore, for $11\leq k\leq n-$ 1, $f_{\text{\rm inv}}$ is $k$th order sum-free if $k\in\{1,n-1\}$ for even $n$ and $k\in\{1,2,n-2,n-1\}$ for odd $n$. Theoretic results and computer searches, which we will summarize in the next section, strongly suggest that these are the only values of $k$ for which $f_{\text{\rm inv}}$ is $k$th order sum-free. This claim has been confirmed in [7] for even $n$, but for odd $n$ the question remains open. (We remind the reader that the even $n$ case does not represent “half” of the question as we will see later that a small prime divisor of $n$ gives an advantage for the treatment of the question.) For $n\geq 2$, define
@@
Because of the last condition, the most difficult cases of Conjecture 1.1 are the scenarios where $n$ is a prime or a product of two primes (not necessarily different) close to each other. Carlet [5] proved that $4\in\mathcal{K}_{n}$ for all $n\geq 6$. We found a different proof of this fact using the Lang-Weil bound. The new proof is included in the present paper because we believe that the approach can be applied to more general situations.
@@
The notion of sum-freedom can be generalized to functions from $\mathbb{F}_{q^{n}}$ to $\mathbb{F}_{q^{n}}$, mutatis mutandis, where $q$ is a prime power. A function $f:\mathbb{F}_{q^{n}}\to\mathbb{F}_{q^{n}}$ is said to be $k$th order sum-free if $\sum_{x\in A}f(x)\neq 0$ for all $k$-dimensional $\mathbb{F}_{q}$-affine subspaces $A$ of $\mathbb{F}_{q^{n}}$. In this context, an immediate question is this: What is the $q$-ary version of the binary multiplicative inverse function $f_{\text{\rm inv}}$ that maintains the properties of $f_{\text{\rm inv}}$? We determine that the “right” $q$-ary version is the function $g_{q-1}:\mathbb{F}_{q^{n}}\to\mathbb{F}_{q^{n}}$ defined by $g_{q-1}(x)=1/x^{q-1}$ (with $1/0$ defined as 00). The reason for this, as we shall see in Section 4, is not entirely obvious. It turns out that most existing results concerning the sum-freedom of $f_{\text{\rm inv}}$ still hold for $g_{q-1}$. However, more interesting things seem to be happening in the $q$-ary case than the binary case. For example, through a computer search, we find that with $q=3,5$ and $n=7$, $g_{q-1}$ on $\mathbb{F}_{q^{7}}$ is $k$th order sum-free for all $1\leq k\leq 6$; similar phenomena have not been observed in the binary case. We are not aware of any immediate applications of the functions $g_{2}(x)=1/x^{2}$ on $\mathbb{F}_{3^{7}}$ and $g_{4}(x)=1/x^{4}$ on $\mathbb{F}_{5^{7}}$. Nevertheless, these are extraordinary examples from a mathematical point of view.
@@
The paper is organized as follows: Section 2 provides the preliminaries for the paper. We recall the existing results about binary sum-free functions, especially the multiplicative inverse function $f_{\text{\rm inv}}$. We also include a determination of all $(n-1)$th order sum-free functions on $\mathbb{F}_{2^{n}}$; this question has not been addressed previously. In Section 3, we confirm Conjecture 1.1 under each of the conditions described above and we prove that $4\in\mathcal{K}_{n}$ for all $n\geq 6$. In Section 4, we explore $q$-ary sum-free functions with focus on the generalization of the binary multiplicative inverse function $f_{\text{\rm inv}}$. It follows from some nontrivial computation that the “right” $q$-ary version of $f_{\text{\rm inv}}$ is the function $g_{q-1}$ defined by $g_{q-1}(x)=1/x^{q-1}$. We show that most existing results concerning the sum-freedom of $f_{\text{\rm inv}}$ still hold for $g_{q-1}$. We also include some limited computer search results on $g_{q-1}$ with $q=3,5$. The appendix contains a proof of the absolute irreducibility of a symmetric homogeneous polynomial of degree 8 in 4 variables over $\mathbb{F}_{2}$. The proof is rather technical and the result is used in Section 3.
@@ 2. Preliminaries
2.1. $k$th order sum-free functions with $k=0,1,2,n-1,n$
@@
  • $f$ is 00th order sum-free if and only if $f(x)\neq 0$ for all $x\in\mathbb{F}_{2^{n}}$;
  • $f$ is $1$st order sum-free if and only if $f$ is a bijection;
  • $f$ is $2$nd order sum-free if and only if $f$ is APN;
  • $f$ is $n$th order sum-free if and only if $\sum_{x\in\mathbb{F}_{2^{n}}}f(x)\neq 0$, i.e., $\deg f<2^{n}-1$ when $f$ is represented by a polynomial modulo $X^{2^{n}}-X$.
There does not appear to be a determination of $(n-1)$th order sum-free functions in the literature. Hence the following theorem is useful for the record.
@@
Then $f$ is $(n-1)$th order sum-free if and only if $a_{2^{n}-1}=0$ and
@@
where $\text{\rm Tr}=\text{\rm Tr}_{2^{n}/2}$. Each hyperplane (affine subspace of co-dimension one) of $\mathbb{F}_{2^{n}}$ is of the form $A_{b}$ or $B_{b}$ (depending on whether it contains 0 or not). For $0\neq b\in\mathbb{F}_{2^{n}}$, we have
@@
By [11, Theorem 2.29], the $\mathbb{F}_{2}$-linear map $\mathbb{F}_{2^{n}}\to\mathbb{F}_{2^{n}}$, $x\mapsto\sum_{j=0}^{n-1}a_{2^{n}-1-2^{j}}x^{2^{j}}$ is bijective if and if (2.1) is satisfied. Therefore, $\sum_{x\in B_{b}}f(x)\neq 0$. Since $a_{2^{n}-1}=0$, we have $\sum_{x\in\mathbb{F}_{2^{n}}}f(x)=0$, hence $\sum_{x\in A_{b}}f(x)=\sum_{x\in B_{b}}f(x)\neq 0$.
@@ 2.2. Known results about $f_{\text{\rm inv}}$
References [4, 5, 7] contain many results concerning the sum-freedom of the multiplicative inverse function $f_{\text{\rm inv}}$ of $\mathbb{F}_{2^{n}}$. We gather them in their strongest versions for the reader’s convenience. We assume that $n\geq 2$ and $11\leq k\leq n-$ 1.
@@
    If $\gcd(k,n)>1$, then $f_{\text{\rm inv}}$ is not $k$th order sum-free (obvious).
@@
    $f_{\text{\rm inv}}$ is $k$th order sum-free if and only if it is $(n-k)$th order sum-free ([5, Corollary 4]).
@@
    If $11\leq k,l\leq n-$ 1 are such that $kl<n$ and $f_{\text{\rm inv}}$ is not $k$th order sum-free nor $l$th order sum-free, then $f_{\text{\rm inv}}$ is not $(k+l)$th order sum-free (update of [5]).
@@
    For $n\geq 6$, $f_{\text{\rm inv}}$ is not $3$rd order sum-free ([5, Corollary 8]).
@@
    For $n\geq 6$, $f_{\text{\rm inv}}$ is not $4$rd order sum-free (update of [5]).
@@
    If $X^{n}-1$ has a factor $X^{k}+a_{k-1}X^{k-1}+\cdots+a_{2}X^{2}+a_{0}\in\mathbb{F}_{2}[X]$, then $f_{\text{\rm inv}}$ is not $k$th order sum-free ([7, Corollary 3.2]).
@@
    If $n\geq 10.8k-5$, then $f_{\text{\rm inv}}$ is not $k$th order sum-free ([7, Theorem 4.2 and Remark 4.4]).

### v2

@@ Abstract.
A function from $\mathbb{F}_{2^{n}}$ to $\mathbb{F}_{2^{n}}$ is said to be $k$th order sum-free if the sum of its values over each $k$-dimensional $\mathbb{F}_{2}$-affine subspace of $\mathbb{F}_{2^{n}}$ is nonzero. This notion was recently introduced by C. Carlet as, among other things, a generalization of APN functions. At the center of this new topic is a conjecture about the sum-freedom of the multiplicative inverse function $f_{\text{\rm inv}}(x)=x^{-1}$ (with $0^{-1}$ defined to be 0). It is known that $f_{\text{\rm inv}}$ is 2nd order (equivalently, (n2)(n-2)th order) sum-free if and only if $n$ is odd, and it is conjectured that for $3\leq k\leq n-3$, $f_{\text{\rm inv}}$ is never $k$th order sum-free. The conjecture has been confirmed for even $n$ but remains open for odd $n$. In the present paper, we show that the conjecture holds under each of the following conditions: (1) $n=13$; (2) $3\mid n$; (3) $5\mid n$; (4) the smallest prime divisor $l$ of $n$ satisfies $(l-1)(l+2)\leq(n+1)/2$. We also determine the “right” $q$-ary generalization of the binary multiplicative inverse function $f_{\text{\rm inv}}$ in the context of sum-freedom. This $q$-ary generalization not only maintains most results for its binary version, but also exhibits some extraordinary phenomena that are not observed in the binary case.
@@ 1. Introduction
In a recent paper [4], Carlet introduced the notion of sum-free functions. A function $f:\mathbb{F}_{2^{n}}\to\mathbb{F}_{2^{n}}$ is said to be $k$th order sum-free if $\sum_{x\in A}f(x)\neq 0$ for all $k$-dimensional affine subspaces $A$ of $\mathbb{F}_{2^{n}}$. Unless specified otherwise, subspaces and affine subspaces of $\mathbb{F}_{2^{n}}$ are over $\mathbb{F}_{2}$. This definition does not involve the multiplicative structure of $\mathbb{F}_{2^{n}}$, hence sum-freedom can be defined for functions from $\mathbb{F}_{2}^{n}$ to $\mathbb{F}_{2}^{n}$. However, the multiplicative structure of $\mathbb{F}_{2^{n}}$ plays an important role in the constructions and analysis of such functions. As described in [4], the motivation for the notion of sum-free functions was two-fold: First, sum-free functions are a natural generalization of almost perfect nonlinear (APN) functions; the latter have been extensively studied in cryptography for their resistance against differential attacks [1, 3, 6, 19, 20, 21]. In particular, APN functions are precisely 2nd order sum-free functions [11]. Second, sum-freedom addresses the vulnerability of block ciphers to integral attacks which exploit the predicability of the sums of values of S-boxes on affine subspaces [14, 22]. Moreover, from a theoretic point of view, sum-freedom is a notion with rich mathematical content; beneath its simple appearance, there are difficult and enticing questions that require sophisticated mathematical tools.
@@
Let $f_{\text{\rm inv}}$ denote the multiplicative inverse function of $\mathbb{F}_{2^{n}}$, that is, $f_{\text{\rm inv}}(x)=x^{-1}$ for $x\in\mathbb{F}_{2^{n}}^{*}$ and $f_{\text{\rm inv}}(0)=0$. For $n\geq 2$, we have $f_{\text{\rm inv}}(x)=x^{2^{n}-2}$ for all $x\in\mathbb{F}_{2^{n}}$. The sum-freedom of $f_{\text{\rm inv}}$ has been the focus of several recent papers [4, 5, 7]. It is well known that $f_{\text{\rm inv}}$ is APN (i.e., 2nd order sum-free) if and only if $n$ is odd [20]. Carlet [4] determined that $\sum_{x\in A}1/x\neq 0$ for all affine subspaces $A$ of $\mathbb{F}_{2^{n}}$ not containing 0. Hence $f_{\text{\rm inv}}$ is $k$th order sum-free if and only if $\sum_{0\neq u\in E}1/u\neq 0$ for all $k$-dimensional subspaces $E$ of $\mathbb{F}_{2^{n}}$. It is trivial that $f_{\text{\rm inv}}$ is 11st order sum-free and it is known from [5] that for $1\leq k\leq n-1$, $f_{\text{\rm inv}}$ is $k$th order sum-free if and only if it is (nk)(n-k)th order sum-free. Therefore, for $1\leq k\leq n-1$, $f_{\text{\rm inv}}$ is $k$th order sum-free if $k\in\{1,n-1\}$ for even $n$ and $k\in\{1,2,n-2,n-1\}$ for odd $n$. Theoretic results and computer searches, which we will summarize in the next section, strongly suggest that these are the only values of $k$ for which $f_{\text{\rm inv}}$ is $k$th order sum-free. This claim has been confirmed in [7] for even $n$, but for odd $n$ the question remains open. (We remind the reader that the even $n$ case does not represent “half” of the question as we will see later that a small prime divisor of $n$ gives an advantage for the treatment of the question.) For $n\geq 2$, define
@@
Because of the last condition, the most difficult cases of Conjecture 1.1 are the scenarios where $n$ is a prime or a product of two primes, not necessarily different, close to each other. Carlet [5] proved that $4\in\mathcal{K}_{n}$ for all $n\geq 6$. We found a different proof of this fact using the Lang-Weil bound. The new proof is included in the present paper because we believe that the approach can be applied to more general situations.
@@
The notion of sum-freedom can be generalized to functions from $\mathbb{F}_{q^{n}}$ to $\mathbb{F}_{q^{n}}$, mutatis mutandis, where $q$ is a prime power. A function $f:\mathbb{F}_{q^{n}}\to\mathbb{F}_{q^{n}}$ is said to be $k$th order sum-free if $\sum_{x\in A}f(x)\neq 0$ for all $k$-dimensional $\mathbb{F}_{q}$-affine subspaces $A$ of $\mathbb{F}_{q^{n}}$. In this context, an immediate question is this: What is the $q$-ary version of the binary multiplicative inverse function $f_{\text{\rm inv}}$ that maintains the properties of $f_{\text{\rm inv}}$? We determine that the “right” $q$-ary version is the function $g_{q-1}:\mathbb{F}_{q^{n}}\to\mathbb{F}_{q^{n}}$ defined by $g_{q-1}(x)=1/x^{q-1}$ with $1/0$ defined as 0. The reason for this, as we shall see in Section 4, is not entirely obvious. It turns out that most existing results concerning the sum-freedom of $f_{\text{\rm inv}}$ still hold for $g_{q-1}$. However, more interesting things seem to be happening in the $q$-ary case than the binary case. For example, through a computer search, we find that with $q=3,5$ and $n=7$, $g_{q-1}$ on $\mathbb{F}_{q^{7}}$ is $k$th order sum-free for all $1\leq k\leq 6$; similar phenomena have not been observed in the binary case. We are not aware of any immediate applications of the functions $g_{2}(x)=1/x^{2}$ on $\mathbb{F}_{3^{7}}$ and $g_{4}(x)=1/x^{4}$ on $\mathbb{F}_{5^{7}}$. Nevertheless, these are extraordinary examples from a mathematical point of view.
@@
The paper is organized as follows: Section 2 provides the preliminaries for the paper. We recall the existing results about binary sum-free functions, especially the multiplicative inverse function $f_{\text{\rm inv}}$. We also include a determination of all (n1)(n-1)th order sum-free functions on $\mathbb{F}_{2^{n}}$; this question has not been addressed previously. In Section 3, we confirm Conjecture 1.1 under each of the conditions described above and we prove that $4\in\mathcal{K}_{n}$ for all $n\geq 6$. In Section 4, we explore $q$-ary sum-free functions with focus on the generalization of the binary multiplicative inverse function $f_{\text{\rm inv}}$. It follows from some nontrivial computation that the “right” $q$-ary version of $f_{\text{\rm inv}}$ is the function $g_{q-1}$ defined by $g_{q-1}(x)=1/x^{q-1}$. We show that most existing results concerning the sum-freedom of $f_{\text{\rm inv}}$ still hold for $g_{q-1}$. We also include some limited computer search results on $g_{q-1}$ with $q=3,5$. The appendix contains a proof of the absolute irreducibility of a symmetric homogeneous polynomial of degree 8 in 4 variables over $\mathbb{F}_{2}$. The proof is rather technical and the result is used in Section 3.
@@ 2. Preliminaries
2.1. $k$th order sum-free functions with k=0,1,2,n1,nk=0,1,2,n-1,n
@@
  • $f$ is 0th order sum-free if and only if $f(x)\neq 0$ for all $x\in\mathbb{F}_{2^{n}}$;
  • $f$ is 11st order sum-free if and only if $f$ is a bijection;
  • $f$ is 22nd order sum-free if and only if $f$ is APN;
  • $f$ is $n$th order sum-free if and only if $\sum_{x\in\mathbb{F}_{2^{n}}}f(x)\neq 0$, i.e., $\deg f=2^{n}-1$ when $f$ is represented by a polynomial modulo $X^{2^{n}}-X$.
The last statement follows from the fact that
$\sum_{x\in\mathbb{F}_{2^{n}}}x^{i}=\begin{cases}0&\text{if}\ 0\leq i\leq 2^{n}-2,\cr 1&\text{if}\ i=2^{n}-1.\end{cases}$
There does not appear to be a determination of (n1)(n-1)th order sum-free functions in the literature. Hence the following theorem is useful for the record.
@@
Then $f$ is (n1)(n-1)th order sum-free if and only if $a_{2^{n}-1}=0$ and
@@
where $\text{\rm Tr}=\text{\rm Tr}_{2^{n}/2}$. Each hyperplane (affine subspace of co-dimension one) of $\mathbb{F}_{2^{n}}$ is of the form $A_{b}$ or $B_{b}$, depending on whether it contains 0 or not. For $0\neq b\in\mathbb{F}_{2^{n}}$, we have
@@
By [13, Theorem 2.29], the $\mathbb{F}_{2}$-linear map $\mathbb{F}_{2^{n}}\to\mathbb{F}_{2^{n}}$, $x\mapsto\sum_{j=0}^{n-1}a_{2^{n}-1-2^{j}}x^{2^{j}}$ is bijective if and if condition (2.1) is satisfied. Therefore, $\sum_{x\in B_{b}}f(x)\neq 0$. Since $a_{2^{n}-1}=0$, we have $\sum_{x\in\mathbb{F}_{2^{n}}}f(x)=0$, hence $\sum_{x\in A_{b}}f(x)=\sum_{x\in B_{b}}f(x)\neq 0$.
@@ 2.2. Known results about $f_{\text{\rm inv}}$
References [4, 5, 7] contain many results concerning the sum-freedom of the multiplicative inverse function $f_{\text{\rm inv}}$ of $\mathbb{F}_{2^{n}}$. We gather them in their strongest versions for the reader’s convenience. We assume that $n\geq 2$ and $1\leq k\leq n-1$.
@@
    If $\gcd(k,n)>1$, then $k\in\mathcal{K}_{n}$ (obvious).
@@
    $f_{\text{\rm inv}}$ is $k$th order sum-free if and only if it is (nk)(n-k)th order sum-free ([5, Theorem 3]).
@@
    If $k,l\in\mathcal{K}_{n}$ and $kl<n$, then $k+l\in\mathcal{K}_{n}$ ([5, Theorem 4]).
@@
    For $n\geq 6$, $3\in\mathcal{K}_{n}$ ([5, Theorem 6]).
@@
    For $n\geq 6$, $4\in\mathcal{K}_{n}$ ([5, Theorem 7]).
@@
    If $X^{n}-1$ has a factor $X^{k}+a_{k-1}X^{k-1}+\cdots+a_{2}X^{2}+a_{0}\in\mathbb{F}_{2}[X]$, then $k\in\mathcal{K}_{n}$ ([7, Corollary 3.2]).
@@
    If $n\geq 13k/3+3$, then $k\in\mathcal{K}_{n}$ ([7, Theorem 4.2]).
2.3. The Moore determinant
The Moore determinant (see [18]) plays an important role in the study of the sum-freedom of $f_{\text{\rm inv}}$ and its $q$-ary version. Let $X_{1},\dots,X_{k}$ be independent indeterminates. The Moore determinant in $X_{1},\dots,X_{k}$ over $\mathbb{F}_{q}$ is defined to be
$\Delta(X_{1},\dots,X_{k})=\left|\begin{matrix}X_{1}&\cdots&X_{k}\cr X_{1}^{q}&\cdots&X_{k}^{q}\cr\vdots&&\vdots\cr X_{1}^{q^{k-1}}&\cdots&X_{k}^{q^{k-1}}\end{matrix}\right|\in\mathbb{F}_{q}[X_{1},\dots,X_{k}].$
Lemma 2.2 ([16, Lemma 3.51]).
We have
Corollary 2.3.
A set of $k$ elements $v_{1},\dots,v_{k}\in\mathbb{F}_{q^{n}}$ are linearly independent over $\mathbb{F}_{q}$ if and only if $\Delta(v_{1},\dots,v_{k})\neq 0$.
For $0\leq i\leq k$, let